# Deferred Commitment Decoding for Diffusion Language Models with Confidence-Aware Sliding Windows

### v1

@@ Abstract
Diffusion language models (DLMs) have recently emerged as a strong alternative to autoregressive models by enabling parallel text generation. To improve inference efficiency and KV-cache compatibility, prior work commonly adopts block-based diffusion, decoding tokens block by block. However, this paradigm suffers from a structural limitation that we term Boundary-Induced Context Truncation (BICT): undecoded tokens near block boundaries are forced to commit without access to nearby future context, even when such context could substantially reduce uncertainty. This limitation degrades decoding confidence and generation quality, especially for tasks requiring precise reasoning, such as mathematical problem solving and code generation. We propose Deferred Commitment Decoding (DCD), a novel, training-free decoding strategy that mitigates this issue. DCD maintains a confidence-aware sliding window over masked tokens, resolving low-uncertainty tokens early while deferring high-uncertainty tokens until sufficient contextual evidence becomes available. This design enables effective bidirectional information flow within the decoding window without sacrificing efficiency. Extensive experiments across multiple diffusion language models, benchmarks, and caching configurations show that DCD improves generation accuracy by 1.39% with comparable time on average compared to fixed block-based diffusion methods, with the most significant improvement reaching 9.0%. These results demonstrate that deferring token commitment based on uncertainty is a simple yet effective principle for improving both the quality and efficiency of diffusion language model decoding.
Deferred Commitment Decoding for Diffusion Language Models with Confidence-Aware Sliding Windows
Yingte Shu Peking University ytshu25@stu.pku.edu.cn          Yuchuan Tian Peking University tianyc@stu.pku.edu.cn          Chao Xu Peking University xuchao@cis.pku.edu.cn
Yunhe Wang Huawei Technologies Co., Ltd yunhe.wang@huawei.com          Hanting Chen Huawei Technologies Co., Ltd chenhanting@huawei.com
Figure 1: An overview of the proposed DCD algorithm.
@@ 1 Introduction
Diffusion language models (DLMs) have recently emerged as a promising alternative to autoregressive models for natural language generation. By decoding tokens in parallel rather than strictly left-to-right, DLMs relax sequential dependencies and enable more flexible generation. Recent models such as NBDiff tian2025next and LLaDA2.0 bie2025llada2 demonstrate that, at comparable scales, DLMs can match or even surpass their autoregressive counterparts on selected reasoning tasks.
@@
A major challenge in practical DLM inference lies in compatibility with key-value (KV) caching. Vanilla DLMs decode tokens in largely unconstrained orders, which prevents effective reuse of cached attention states and leads to slow inference. To address this issue, block-based diffusion methods have been proposed arriola2025block, partitioning the sequence into blocks that are decoded sequentially while allowing parallel decoding within each block. This semi-autoregressive structure significantly improves KV-cache efficiency and has become a standard design choice in recent DLM systems.
@@
Despite their efficiency benefits, block-based diffusion methods introduce a fundamental limitation, which we refer to as Boundary-Induced Context Truncation (BICT). Once decoding proceeds to the next block, undecoded tokens in the current block are forced to commit, even if nearby future tokens—often only a few positions away—could provide crucial disambiguating context. This issue is particularly harmful for tokens in semantically critical positions, where insufficient context leads to low-confidence decisions and error propagation. Importantly, this limitation is not caused by an incorrect decoding order, but by rigid block boundaries that assume information sufficiency at block completion.
@@
Our core hypothesis is that decoding quality can be improved by deferring commitment on high-uncertainty tokens until sufficient contextual evidence becomes available, without abandoning the efficiency advantages of block-based decoding. Based on this insight, we propose Deferred Commitment Decoding (DCD), a training-free decoding strategy that replaces fixed block boundaries with a confidence-aware sliding window. Within this window, tokens with low uncertainty are resolved first, while high-uncertainty tokens remain masked and continue to benefit from dynamically expanding context. This mechanism enables localized bidirectional information flow while preserving compatibility with existing caching schemes.
@@
We evaluate DCD on a diverse set of tasks, including mathematical reasoning math500; gsm8k, code generation mbpp; humaneval, and instruction following ifeval, using multiple diffusion language models nie2025large; ye2025dream; wu2025fast2 and various KV caching configurations. Across all settings, DCD consistently improves generation accuracy by 1.39% with comparable time on average compared to fixed block-based diffusion baselines, while the maximum improvement in certain configurations reaches 9.0%. These results establish DCD as a strong state-of-the-art decoding method for DLMs.
@@
  • We demonstrate that DCD achieves consistent accuracy improvements and decoding speedups over fixed block-based diffusion methods across models, tasks, and caching configurations.
@@ 2.1 DLMs Taxonomy
There are two main lines of work that adapt diffusion techniques ho2020denoising from computer vision to natural language processing. Continuous diffusion language models li2022diffusion; gong2022diffuseq project discrete language tokens into continuous spaces and apply denoising processes to recover text outputs. In contrast, discrete diffusion language models draw inspiration from masked language modeling devlin2019bert, gradually recovering masked tokens at predefined generation slots. Compared to continuous approaches, discrete DLMs better align with the inherently discrete nature of language and can be more easily adapted from existing autoregressive models gong2024scaling; as a result, they have become the dominant paradigm in recent diffusion-based language modeling research. Unless otherwise specified, we use the term DLMs in this paper to refer to discrete diffusion language models.
Discrete DLMs typically employ one of two attention mechanisms in the Transformer architecture: semi-causal attention or full attention. Models such as BD3-LM arriola2025block and Fast-dLLMv2 wu2025fast2 adopt blockwise causal attention, where tokens attend only to their current block and preceding blocks. In contrast, models such as LLaDA nie2025large and Dream ye2025dream adopt full attention, allowing each token to condition on the entire sequence during decoding. In this work, we consider both semi-causal and full-attention DLMs to demonstrate the generality and robustness of the proposed DCD decoding algorithm across different architectural choices.
2.2 Use of Cache in DLMs
Key-value (KV) caching is a crucial optimization for Transformer-based diffusion language models, as it enables faster inference by reusing previously computed attention states. For full-attention DLMs, Fast-dLLM wu2025fast enforces a blockwise decoding order and introduces prefix and dual caching strategies, where key-value pairs are cached for prefix blocks and for both prefix and suffix blocks, respectively. dKV-cache ma2025dkv proposes a delayed KV-cache mechanism, in which tokens are cached one step after being decoded. A further variant selectively caches only the neighboring tokens of the most recently decoded positions, thereby reducing cache inconsistency and computational overhead.
Due to the bidirectional nature of full-attention DLMs, KV caching in these models is inherently approximate, necessitating periodic cache refreshes to maintain correctness. In contrast, semi-causal attention DLMs naturally support exact prefix KV caching without refresh, as later decoding steps do not alter the attention context of earlier tokens. Fast-dLLMv2 wu2025fast2 further extends this design by introducing intra-block dual caching, where both prefix and suffix “sub-block” tokens within a block are cached to improve efficiency during blockwise decoding.
Overall, these caching strategies substantially accelerate DLM inference, although they may introduce minor degradation in generation quality due to cache approximation. The proposed DCD algorithm is compatible with existing KV caching schemes in both full-attention and semi-causal DLMs, allowing it to achieve efficiency gains without sacrificing decoding accuracy.
@@ 2.3 Decoding Strategies of DLMs
Earlier works on DLMs austin2021structured; sahoo2024simple randomly unmask and remask a fixed number of tokens at each decoding step, which often yields suboptimal performance. Later approaches incorporate confidence- or entropy-based criteria, decoding tokens whose uncertainty exceeds a threshold or falls within Top-$k$ candidates. These strategies improve flexibility and parallelism but still rely on fixed decoding ranges.
[…]
@@ 3.2 Block-based decoding of DLMs
where $\mathcal{V}$ is the vocabulary set and $\mathcal{S}^{(t)}=\{i\in\mathcal{M}^{(t)}\mid\bigwedge_{t}\text{Cond}_{t}(i)\}$ specifies whether position $i$ is eligible for decoding at the current step.
In block-based decoding, the eligibility condition constrains decoding positions to a fixed region, typically the current block: $\text{Cond}_{0}(i)=\bigl[i\in\mathcal{B}_{\text{cur}}\bigr]$. Full-attention DLMs may optionally adopt block-based decoding to improve KV-cache compatibility. In contrast, semi-causal DLMs must employ block-based decoding due to their blockwise attention constraints. Within a large attention block, semi-causal DLMs may further apply sub-block decoding, where decoding positions are restricted to a smaller contiguous region: $\text{Cond}_{0}(i)=\bigl[i\in\mathcal{B}_{\text{cur}_{1}:\text{cur}_{2}}\bigr]$, where $\mathcal{B}_{\text{cur}_{1}:\text{cur}_{2}}$ denotes a contiguous subrange of blocks within a larger attention block.
@@ 4 Boundary-Induced Context Truncation
The advantages of DLMs over their autoregressive counterparts lie primarily in their (semi-)bidirectional attention horizons. piskorz2025masks found that DLMs exhibit a strong contextual locality bias, in which nearby tokens contribute disproportionately to prediction confidence. Informally, this can be expressed as:
@@
where $b<i+\omega_{r}$ denotes the right boundary of the current block. We term this reduction in a token’s effective contextual window the Boundary-Induced Context Truncation phenomenon. Although these tokens receive insufficient context, they must be decoded before entering the next block under the block-based paradigm. Consequently, this leads to low-confidence decoding at the end of each block and ultimately degrades the generation performance of DLMs.
Will increasing the block size solve this problem? Increasing the block size partially reduces the number of blocks, but it does not break the rigidity of block boundaries. Furthermore, it may lead to the long decoding-window problem seo2025fast and weaken the effectiveness of KV caching. As additionally evidenced by previous experiments wu2025fast, setting an excessively large block size is therefore not a desirable solution.
@@ 5 Deferred Commitment Decoding
5.1 Three Rules of the DCD Algorithm
@@
  • (1)
    Identify tokens that suffer from the BICT phenomenon;
@@
As illustrated in Figure 1, the proposed Deferred Commitment Decoding (DCD) algorithm maintains a sliding window and defers the decoding of low-confidence tokens. It follows three rules to achieve the above goals:
Rule (1): Decoding is based on confidence criteria within the window. The prediction confidence of masked tokens serves as a key indicator for identifying BICT-affected tokens, as low confidence strongly suggests insufficient context. Similar to prior work, we set a confidence threshold $\tau_{\text{conf}}$, and only masked tokens within the decoding window whose confidence exceeds $\tau_{\text{conf}}$ are decoded. Specifically, if none of the tokens exceeds the threshold, the most confident token is decoded.
Rule (2): The decoding window slides left to right with constraints. The sliding window defines the range of tokens eligible for decoding. It abandons fixed boundaries across consecutive decoding steps; instead, it moves from left to right within the generation slot of the DLM. Formally, let $L^{(t)},R^{(t)})$ denote the left and right endpoints of the sliding window, and let $\mathbf{x}^{(t)}_{[l:r]}$ denote the generation slot at decoding step $t$. Then,
@@
Rule (3): Tokens are deferred until they are sufficiently confident. In block-based decoding, low-confidence tokens may be forcibly decoded to complete a block. In contrast, the proposed DCD algorithm handles low-confidence tokens more gracefully: it defers their decoding and expands the available context until the DLM becomes sufficiently confident to predict them. This approach significantly reduces low-confidence decoding events at later stages, thereby improving overall performance.
How does the DCD algorithm differ from AdaBlock? The AdaBlock method lu2025adablock employs adaptive block sizes based on delimiter semantics in the generated tokens, improving generation coherence. However, once computed, the block sizes remain fixed; as a result, AdaBlock may still suffer from BICT and thus leaves room for further improvement.
@@
0: Generation slot $\mathbf{x}^{(t)}_{[l:r]}$, DLM model $p_{\theta}(\cdot\mid\mathbf{x}^{(t)})$, window parameters $s_{\text{init}},s_{\text{max}}$, cache parameters $\text{cache\_type},B^{\prime},r$, confidence threshold $\tau_{\text{conf}}$.
1: Initialize $L^{(t)},R^{(t)}=l,l+s_{\text{init}}$.
2: Initialize cache refresh countdown $cd=0$.
3:while $\mathcal{M}^{(t)}\neq\emptyset$ do
4:  Obtain eligible masked tokens $\mathcal{E}^{(t)}=[L^{(t)},R^{(t)})\cap\mathcal{M}^{(t)}$.
5:  if cache_type $\neq$ none and $cd\leq 0$ then
6:   Refresh the cache based on cache_type and Equation 8.
7:   Set $cd=B^{\prime}$.
8:  end if
9:  For all $i\in\mathcal{E}^{(t)}$, compute confidence $c_{i}=\max_{x_{i}\in\mathcal{V}}p_{\theta}(x_{i}\mid\mathbf{x}^{(t)})$.
10:  Select decoding positions $\mathcal{S}^{(t)}=\{i\mid c_{i}\geq\tau_{\text{conf}}\}\cup\arg\max_{i}\{c_{i}\}$.
11:  Update $\mathbf{x}^{(t-1)}$ with $\mathcal{S}^{(t)}$ using Equation 3.
12:  Update $L^{(t-1)},R^{(t-1)}$ with $\mathbf{x}^{(t-1)}_{[l:r]}$ using Equations 6 and 7.
13:  Update $cd=cd-|\mathcal{S}^{(t)}|$.
14:  Update t=t1t=t-1.
15:end while
16:return Final sequence $\mathbf{x}^{(t)}_{[l:r]}$.
5.2 Applying DCD to Different Types of DLMs
The generation slot of DCD is aligned with the bidirectional attention intervals of different types of DLMs.
For full-attention DLMs, there is a single generation slot consisting of the prefilled masked tokens following the prompt tokens, which represents the fixed-length output of the corresponding query. In this case, the DCD algorithm completely replaces block-based decoding.
For semi-causal DLMs, multiple generation slots arise during inference, each corresponding to a “large block” predefined by the attention structure. In this setting, the training-free DCD algorithm cannot modify the blockwise decoding pattern at the macro level, as these DLMs are trained with fixed block sizes. However, DCD can replace fixed-length sub-block decoding within each large block and outperform both standard block-based and sub-block-based decoding methods.

### v2

@@ Abstract
Diffusion language models (DLMs) have recently emerged as a strong alternative to autoregressive models by enabling parallel text generation. To improve inference efficiency and KV-cache compatibility, prior work commonly adopts block-based diffusion, decoding tokens block by block. However, this paradigm suffers from a structural limitation that we term Boundary-Induced Context Truncation (BICT): undecoded tokens near block boundaries are forced to commit without access to nearby future context, even when such context could substantially reduce uncertainty. This limitation degrades decoding certainty and generation quality, especially for tasks requiring precise reasoning, such as mathematical problem solving and code generation. We propose Deferred Commitment Decoding (DCD), a novel, training-free decoding strategy that mitigates this issue. DCD maintains a certainty-aware sliding window over masked tokens, resolving low-uncertainty tokens early while deferring high-uncertainty tokens until sufficient contextual evidence becomes available. Extensive experiments across multiple diffusion language models, benchmarks, and caching configurations show that DCD improves generation accuracy by 1.73% with comparable time on average compared to fixed block-based diffusion methods, with the most significant improvement reaching 16.5%. These results demonstrate that deferring token commitment based on uncertainty is a simple yet effective principle for improving both the quality and efficiency of diffusion language model decoding. Code: https://github.com/shuyingte/DCD
Figure 1: An overview of the proposed DCD algorithm. DCD defers uncertain tokens’ decoding with sliding windows for DLMs, thereby improving performance.
@@ 1 Introduction
Diffusion language models (DLMs) have recently emerged as a promising alternative to autoregressive models for natural language generation. By decoding tokens in parallel rather than strictly left-to-right, DLMs relax sequential dependencies and enable more flexible generation. Recent models such as NBDiff Tian et al. (2025) and LLaDA2.0 Bie et al. (2025) demonstrate that, at comparable scales, DLMs can match their autoregressive counterparts on selected reasoning tasks.
@@
A major challenge in practical DLM inference lies in compatibility with key-value (KV) caching. Vanilla DLMs decode tokens in largely unconstrained orders, which prevents effective reuse of cached attention states and leads to slow inference. To address this issue, block-based diffusion methods have been proposed Arriola et al. (2025); Nie et al. (2025), partitioning the sequence into blocks that are decoded sequentially while allowing parallel decoding within each block. This semi-autoregressive structure significantly improves KV-cache efficiency and has become a standard design choice in recent DLM systems.
@@
Despite their efficiency benefits, block-based diffusion methods introduce a fundamental limitation, which we refer to as Boundary-Induced Context Truncation (BICT). Once decoding proceeds to the next block, undecoded tokens in the current block are forced to commit, even if nearby future tokens—often only a few positions away—could provide crucial disambiguating context. This issue is particularly detrimental for tokens in semantically critical positions, where insufficient context leads to low-certainty decisions and error propagation. Importantly, this limitation is not caused by an incorrect decoding order but by rigid block boundaries that assume information sufficiency upon block completion.
@@
Our core hypothesis is that decoding quality can be improved by deferring commitment on high-uncertainty tokens until sufficient contextual evidence becomes available, without abandoning the efficiency advantages of block-based decoding. Based on this insight, we propose Deferred Commitment Decoding (DCD), a training-free decoding strategy that replaces fixed block boundaries with a certainty-aware sliding window. Within this window, tokens with low uncertainty are resolved first, while high-uncertainty tokens remain masked and continue to benefit from dynamically expanding context. This mechanism enables localized bidirectional information flow while preserving compatibility with existing caching schemes.
@@
We evaluate DCD on a diverse set of tasks, including mathematical reasoning Lightman et al. (2023); Cobbe et al. (2021), code generation Austin et al. (2021b); Chen et al. (2021), and instruction following Zhou et al. (2023), using multiple diffusion language models Nie et al. (2025); Ye et al. (2025); Wu et al. (2025a); Tian et al. (2025) and various KV caching configurations. Across all settings, DCD consistently improves generation accuracy by +1.73% with comparable inference time on average compared to fixed block-based diffusion baselines, with the maximum improvement in certain configurations reaching +16.5%. These results establish DCD as a strong state-of-the-art decoding method for DLMs.
@@
  • We demonstrate that DCD achieves consistent accuracy improvements over fixed block-based diffusion methods across models, tasks, and caching configurations.
@@ 2.1 DLMs Taxonomy
There are two main lines of work that adapt diffusion techniques Ho et al. (2020) from computer vision to natural language processing. Continuous diffusion language models Li et al. (2022); Gong et al. (2022) project discrete language tokens into continuous spaces and apply denoising processes to recover text outputs. In contrast, discrete diffusion language models draw inspiration from masked language modeling Devlin et al. (2019), gradually recovering masked tokens at predefined generation slots. Compared to continuous approaches, discrete DLMs better align with the inherently discrete nature of language and can be more easily adapted from existing autoregressive models Gong et al. (2024); as a result, they have become the dominant paradigm in recent diffusion-based language modeling research. Unless otherwise specified, we use the term DLMs in this paper to refer to discrete diffusion language models.
Discrete DLMs typically employ one of two attention mechanisms in the Transformer architecture: semi-causal attention or full attention. Models such as BD3-LM Arriola et al. (2025), NBDiff Tian et al. (2025), and Fast-dLLMv2 Wu et al. (2025a) adopt semi-causal attention, which indicates that their predefined attention maps are incomplete and tokens attend only to their current ”block” and previous ones. In contrast, models such as LLaDA Nie et al. (2025) and Dream Ye et al. (2025) adopt full attention, allowing each token to condition on the entire sequence during decoding. In this work, we consider both semi-causal and full-attention DLMs to demonstrate the generalization and robustness of the proposed DCD decoding algorithm across different architectural choices.
@@ 2.2 Decoding Strategies of DLMs
Earlier works on DLMs Austin et al. (2021a); Sahoo et al. (2024) randomly unmask and remask a fixed number of tokens at each decoding step, which often yields suboptimal performance. Later approaches incorporate confidence- or entropy-based criteria, decoding tokens whose uncertainty falls below a threshold or lies within the top-k candidates. These strategies improve flexibility and parallelism but still rely on fixed decoding ranges.
More recently, a variety of decoding strategies have been introduced to enhance either the performance or computational efficiency of discrete diffusion language models (DLMs). Among training-based approaches, Xu et al. (2024) employs energy functions to steer the decoding process, yielding a 1.3x speedup alongside notable gains in generation quality. FS-DFM Monsefi et al. (2025) formulates a discrete flow-matching framework that generates 1024 tokens in just eight sampling steps without degrading perplexity, while SDLM Liu et al. (2025) adaptively decodes token sequences based on prediction confidence.
For training-free methods, Fu et al. (2025) introduces an Explore-Then-Exploit scheduling mechanism that maximizes information gain per decoding round to improve efficiency. Chen et al. (2025) enhances decoding quality by leveraging historical trajectories to inform current predictions, and Li et al. (2025) proposes early-commit decoding to accelerate inference in DLMs while largely preserving output fidelity. Despite these advances, existing training-free strategies still lack mechanisms to dynamically adjust the decoding horizon or to enrich contextual support for low-certainty tokens, indicating significant opportunities for further innovation.
@@ 3.2 Block-based decoding of DLMs
where $\mathcal{V}$ is the vocabulary set and $\mathcal{S}^{(t)}\subseteq\mathcal{M}^{(t)}$ specifies the positions eligible for decoding at the current step.
In block-based decoding, the eligibility condition constrains decoding positions to a fixed region, typically the current block: $\mathcal{S}^{(t)}\subseteq\{i\mid i\in\mathcal{B}_{\text{cur}}\}$. Full-attention DLMs may optionally adopt block-based decoding to improve KV-cache compatibility. In contrast, semi-causal DLMs must employ block-based decoding due to their blockwise attention constraints. Within a large attention block, semi-causal DLMs may further apply sub-block decoding, where decoding positions are restricted to a smaller contiguous region: $\mathcal{S}^{(t)}\subseteq\{i\mid i\in\mathcal{B}_{\text{cur}_{1}:\text{cur}_{2}}\}$, where $\mathcal{B}_{\text{cur}_{1}:\text{cur}_{2}}$ denotes a contiguous subrange of blocks within a larger attention block.
@@ 4 Boundary-Induced Context Truncation
The advantages of DLMs over their autoregressive counterparts lie primarily in their (semi-)bidirectional attention horizons. Piskorz et al. (2025) found that DLMs exhibit a strong contextual locality bias, in which nearby tokens $\mathbf{x}_{[i-\omega_{l}:i+\omega_{r}]}$ contribute disproportionately to prediction certainty of $i$-th token. Informally, this can be expressed as:
@@
where $b<i+\omega_{r}$ denotes the right boundary of the current block. We term this reduction in a token’s effective contextual window the Boundary-Induced Context Truncation phenomenon. Although these tokens receive insufficient context, they must be decoded before proceeding to the next block under the block-based paradigm. Consequently, this leads to low-certainty decoding at the end of each block and ultimately degrades the generation performance of DLMs.
Figure 2: An example of DCD and block-based decoding. DCD outperforms block-based decoding by solving BICT problem. More details can be found in Appendix B due to page limit.
Figure 3: Confidence distributions of various models. We use LLaDA-8B-Instruct, Dream-v0-Base-7B and Fast-dLLM-v2-7B on GSM8K with DCD and (sub-)block-based decoding using dual cache and log10 scale for clarity. The DCD algorithm yields fewer low-certainty decoding steps across these models.
@@ 5 Deferred Commitment Decoding
5.1 Core Design of the DCD Algorithm
@@
As illustrated in Figure 1, the proposed Deferred Commitment Decoding (DCD) algorithm maintains a sliding window and defers the decoding of low-certainty tokens. It follows two principal design principles to achieve the above goals:
Design (1): The decoding window slides left to right with constraints. The sliding window defines the range of tokens eligible for decoding. It abandons fixed boundaries across consecutive decoding steps; instead, it moves from left to right within the generation slot of the DLM. Formally, let $L^{(t)},R^{(t)})$ denote the left and right endpoints of the sliding window, and let $\mathbf{x}^{(t)}_{[l:r]}$ denote the generation slot at decoding step $t$. Then,
@@
Design (2): Tokens are deferred from decoding until sufficiently certain given nearby context. Low prediction certainty—measured by the confidence or entropy of masked tokens—serves as a key indicator for identifying BICT-affected tokens with insufficient context. In block-based decoding, tokens may be forcibly decoded to complete a block. In contrast, the proposed DCD algorithm handles them more gracefully: masked tokens are decoded only when their certainty exceeds threshold $\tau$ or ranks among the top in the window.
$\mathcal{S}^{(t)}=\mathcal{M}^{(t)}\cap[L^{(t)},R^{(t)})\cap\{i\mid\mathcal{C}(i)\geq\min(\tau,\max_{j}\mathcal{C}(j))\}$ (8)
where the certainty metric $\mathcal{C}(i)$ of the $i$-th token is computed using confidence or negative entropy. This approach significantly reduces low-certainty decoding events at later stages, thereby improving overall performance.
How does the DCD algorithm differ from AdaBlock? The AdaBlock method Lu et al. (2025) employs adaptive block sizes based on delimiter semantics in the generated tokens, improving generation coherence. However, once determined, the block sizes remain fixed; as a result, AdaBlock may still suffer from BICT and thus leaves room for further improvement.
@@
0: Generation slot $\mathbf{x}^{(t)}_{[l:r]}$, DLM $p_{\theta}(\cdot)$, window parameters $s_{\text{init}},s_{\text{max}}$, cache parameters $\text{cache\_type},B^{\prime},r$, certainty threshold $\tau$, DBE parameters $\tau_{\text{low}},e_{\text{step}},e_{\text{max}}$.
1: Initialize $L^{(t)},R^{(t)}=l,l+s_{\text{init}}$, cache refresh $cd=0$.
2:while $\mathcal{M}^{(t)}\neq\emptyset$ do
3:  if cache_type $\neq$ none and $cd\leq 0$ then
4:   Refresh the cache based on Equation 11.
5:   Set $cd=B^{\prime}$.
6:  end if
7:  if Equation 9 is hold then
8:   Update $r,R^{(t-1)}$ based on Equation 10,7.
9:   continue
10:  end if
11:  Select decoding positions $\mathcal{S}^{(t)}$ using Equation 8.
12:  Update $\mathbf{x}^{(t-1)}$ with $\mathcal{S}^{(t)}$ using Equation 3.
13:  Update $L^{(t-1)},R^{(t-1)}$ with $\mathbf{x}^{(t-1)}_{[l:r]}$ using Equations 6 and 7.
14:  Update $cd=cd-|\mathcal{S}^{(t)}|$.
15:  Update t=t1t=t-1.
16:end while
17:return Final sequence $\mathbf{x}^{(t)}_{[l:r]}$.
5.2 Dynamic Block Extension for Semi-causal DLMs
The core design of DCD fits well with full-attention DLMs because their generation slots span the entire sequence, and DCD completely replaces block-based decoding. However, for semi-causal DLMs, the large block size is predefined, and vanilla DCD operates only at the intra-block level by replacing fixed-length sub-block decoding with small sliding windows. Therefore, we propose Dynamic Block Extension (DBE), a patch to DCD for semi-causal DLMs trained with multiple block sizes. Specifically, when a low-certainty token is about to be committed and the window’s sliding is blocked by the right boundary of the large block—i.e., the BICT problem arises due to rigid block boundaries—the following condition holds:
$\begin{split}\min_{i\in\mathcal{S}^{(t)}}\mathcal{C}(i)<\tau_{\text{low}}\quad\text{and}\quad R^{(t)}-L^{(t)}<s_{\text{max}}\\ \text{and}\quad\sum_{k=L^{(t)}}^{R^{(t)}-1}\left[x^{(t)}_{k}=\langle\text{MASK}\rangle\right]<s_{\text{init}}.\end{split}$ (9)
When this occurs, DBE aborts the current decoding step and expands the block with an upper limit:
$r^{\prime}=\min(r+e_{\text{step}},\ \text{blocksize}+e_{\text{max}}).$ (10)
Afterward, the sliding window is recalculated according to Equation 7, and a new decoding step begins to continue the generation loop. Although DBE relies on the DLM’s ability to generalize to variable block sizes, this patch can further enhance performance for models appropriately trained for such flexibility.